OpenStudy (anonymous):

Help Please! Chad bought a new graduated cylinder for his chemistry class. It holds 305.76 milliliters of liquid. If the cylinder has a radius of 2.8 cm, then how tall is the cylinder? (1 milliliter is equal to 1 cubic centimeter) A.54.6 cm B.109.2 cm C.11,924.6400000000 cm D.39 cm

OpenStudy (anonymous):

the formula for a cylinder is what? try finding out

OpenStudy (anonymous):

12.41...

OpenStudy (anonymous):

\[pie*r*h\]

OpenStudy (anonymous):

305.76=3.14*2.8*h

OpenStudy (anonymous):

will that get me the height?

OpenStudy (anonymous):

305.76=8.792*h

OpenStudy (anonymous):

305.76/ 8.792= h

OpenStudy (anonymous):

34.777

OpenStudy (anonymous):

close to d

OpenStudy (anonymous):

no bandit the volume =pi* r^2 * h

OpenStudy (anonymous):

oh.....

OpenStudy (anonymous):

well.. sorry then

OpenStudy (anonymous):

but no choice is true ???

OpenStudy (anonymous):

and this is why i don't like study island...